OpenStudy (anonymous):

Integrate from 1 to 4 (5x^2-x)/(2(x)^(1/2))

hartnn (hartnn):

hint: ,separate the denominator. (5x^2-x)/(2(x)^(1/2)) = 5x^2/(2(x)^(1/2)) - x / (2(x)^(1/2))

OpenStudy (anonymous):

Would u do u sub after that or bring the power to the top (making it negative)

hartnn (hartnn):

after simplification, power(exponent) won't become negative.... use this \(\huge \frac{x^m}{x^n}=x^{m-n}\)

hartnn (hartnn):

for first term, \(\huge \frac{x^2}{x^{1/2}}=x^{2-1/2}\)

OpenStudy (anonymous):

(